A neutron of mass m collides elastically with a nucleus of mass M which is at rest .if the initial kinetic energy of neutron is K ,Calculate the maximum kinetic energy that it can lose during the collision :
1
Expert's answer
2018-10-24T11:03:09-0400

ΔK/K=1-((M-m)/(M+m))^2

ΔK/K=1-((1-m/M)/(1+m/M))^2

The maximum kinetic energy that it can lose during the collision:

ΔK=K(1-((1-m/M)/(1+m/M))^2 )